Bill is twice as old as his brother Dan. In 7 years, Bill will be only one and one-half times as old as Dan. How old is Bill now?

Answer by stanbon(75887)

Bill is twice as old as his brother Dan. In 7 years, Bill will be only one and one-half times as old as Dan. How old is Bill now?
----------------------
Let Bill's age NOW be "B".
Let Dan's age NOW be "D".
---
Equations:
B = 2D
(B+7)= (3/2)(D+7)
-------------------------
Substitute for B and solve for "D":
(2D+7) = (3/2)(D+7)
Multiply both sides by 2 to get:
4D+14 = 3D + 21
D = 7 (Dan's age NOW)
---
Since B = 2D, B = 14 (Bill's age NOW)
=========================================
Cheers,
Stan H.